# Greedy Algorithm for Structured Bandits: A Sharp Characterization of Asymptotic Success / Failure*

### v1

@@ Appendix D $\mathtt{StructuredMAB}$ Failure: deterministic decoy arm
Proof E.2.
The proof is similar to that of Lemma B.1 in the previous section, which is a Hoeffding-style concentration bound with a union bound. We can simply treat each context-arm pair $(x,a)$ as an arm, and this directly yields the result.
@@
Lemma E.3.
Assume the event in Lemma E.1 holds. For any round $s\in[t]$, the cumulative loss at the true underlying function, $\mathtt{MSE}_{s}(f^{*})$, can be upper bounded as $|\mathcal{X}|K\cdot O(\log t)$.
Proof E.4.
@@
Lemma E.5.
Assume the event in Lemma E.1 holds. Fix any round $s$. Let $T^{\prime}$ be some suitably chosen parameter and $T^{\prime}=|\mathcal{X}|K/\Gamma(f^{*})^{2}\cdot O(\log(|\mathcal{X}|Kt))$. Suppose $\mathtt{Greedy}$ executes some suboptimal policy $\pi$ in round $s$. Then there exists context $x$, such that $N_{s}(x,\pi(x))<T^{\prime}$.
Proof E.6.
@@
Hence, any function satisfying eq. E.1 cannot possibly minimize $\mathtt{MSE}_{s}(\cdot)$. In other words, the function minimizing the loss at this step $f_{t}$ must satisfy
@@
Lemma E.7.
Conditional on the event in Lemma E.1, the expected total number of times of suboptimal policy execution is no larger than |$\absolutevalue{\mathcal{X}}KT^{\prime}/p_{0}=(\absolutevalue{\mathcal{X}}K/\Gamma(f^{*}))^{2}/p_{0}\cdot O(\log t)$.
Proof E.8.
@@
Appendix F $\mathtt{StructuredCB}$ Failure: Proof of Theorem 4.3(b)
@@
Appendix G Proofs for $\mathtt{DMSO}$
@@
G.1 $\mathtt{DMSO}$ Success: Proof of Theorem 5.3(a)
@@
G.2 $\mathtt{DMSO}$ Failure: Proof of Theorem 5.3(b)
@@
Proof G.8.
Fix some model $M$. If the distribution for $M(\pi_{\mathtt{dec}})$ and $M_{\mathtt{dec}}(\pi_{\mathtt{dec}})$ were the same, then $\Psi_{j}$ would be 0 for any $j$. Hence we can assume $M(\pi_{\mathtt{dec}})$ and $M_{\mathtt{dec}}(\pi_{\mathtt{dec}})$ are two different distributions. Then $\Psi_{j}$ would be a subgaussian random variable with $\operatornamewithlimits{\mathbb{E}}[\Psi_{j}]>\Gamma(M_{\mathtt{dec}})$. By the previous Lemma G.3, the event $E_{2}$ holds specifically for model $M$ with probability at least 1 - $\delta/|M|$. Then, by a union bound, event $E_{2}$ holds with probability $1-\delta$.
Lemma G.9.

### v2

@@ Appendix B Novelty of self-identifiability
Proof D.2.
The proof is similar to that of Lemma C.1 in the previous section, which is a Hoeffding-style concentration bound with a union bound. We can simply treat each context-arm pair $(x,a)$ as an arm, and this directly yields the result.
@@
Lemma D.3.
Assume the event in Lemma D.1 holds. For any round $s\in[t]$, the cumulative loss at the true underlying function, $\mathtt{MSE}_{s}(f^{*})$, can be upper bounded as $|\mathcal{X}|K\cdot O(\log t)$.
Proof D.4.
@@
Lemma D.5.
Assume the event in Lemma D.1 holds. Fix any round $s$. Let $T^{\prime}$ be some suitably chosen parameter and $T^{\prime}=|\mathcal{X}|K/\Gamma(f^{*})^{2}\cdot O(\log(|\mathcal{X}|Kt))$. Suppose $\mathtt{Greedy}$ executes some suboptimal policy $\pi$ in round $s$. Then there exists context $x$, such that $N_{s}(x,\pi(x))<T^{\prime}$.
Proof D.6.
@@
[…]
@@
Lemma D.7.
Conditional on the event in Lemma D.1, the expected total number of times of suboptimal policy execution is no larger than |$\absolutevalue{\mathcal{X}}KT^{\prime}/p_{0}=(\absolutevalue{\mathcal{X}}K/\Gamma(f^{*}))^{2}/p_{0}\cdot O(\log t)$.
Proof D.8.
@@
D.2 $\mathtt{StructuredCB}$ Failure: Proof of Theorem 4.3(b)
@@
Appendix E $\mathtt{DMSO}$ characterization: Proof of Theorem 5.3
@@
E.1 $\mathtt{DMSO}$ Success: Proof of Theorem 5.3(a)
@@
E.2 $\mathtt{DMSO}$ Failure: Proof of Theorem 5.3(b)
@@
Proof E.8.
Fix some model $M$. If the distribution for $M(\pi_{\mathtt{dec}})$ and $M_{\mathtt{dec}}(\pi_{\mathtt{dec}})$ were the same, then $\Psi_{j}$ would be 0 for any $j$. Hence we can assume $M(\pi_{\mathtt{dec}})$ and $M_{\mathtt{dec}}(\pi_{\mathtt{dec}})$ are two different distributions. Then $\Psi_{j}$ would be a subgaussian random variable with $\operatornamewithlimits{\mathbb{E}}[\Psi_{j}]>\Gamma(M_{\mathtt{dec}})$. By the previous Lemma E.3, the event $E_{2}$ holds specifically for model $M$ with probability at least 1 - $\delta/|M|$. Then, by a union bound, event $E_{2}$ holds with probability $1-\delta$.
Lemma E.9.
@@
Appendix F $\mathtt{StructuredMAB}$ with an Infinite Function Class: Proof of Theorem 6.2
F.1 Success: Proof of Theorem 6.2(a)
The proof of Theorem 3.3(a) carries over with a modified version of Lemma C.5. Thus, it suffices to state and prove this modified lemma.
Lemma F.1.
Let $\mathcal{F}$ be an infinite function class. Assume the event in Lemma C.1. The number of times any suboptimal arm is chosen cannot exceed $T^{\prime}$ rounds, where $T^{\prime}$ is some parameter with $T^{\prime}=(K/\varepsilon^{2})\cdot O(\log t)$.
Proof F.2.
We prove this by contradiction. Consider any round $\tau$ during which some suboptimal arm $a$ is chosen above this threshold $T^{\prime}$. The reward for arm $a$ is going to get concentrated within $O(\varepsilon)$ to $f^{*}(a)$, in particular:
|$\absolutevalue{\bar{r}_{\tau}(a)-f^{*}(a)}<\varepsilon/2.$
Take any reward vector $f^{\prime}$ such that |$\absolutevalue{f(a^{\prime})-f(a)}>\varepsilon$. Then we have
|$\absolutevalue{\bar{r}_{\tau}(a)-f(a)}\geq\varepsilon/2.$
Then the cumulative loss
$\mathtt{MSE}_{\tau}(f^{\prime})\geq T^{\prime}\cdot(\varepsilon/2)^{2}=K\cdot\Omega(\log t).$
Therefore any $f^{\prime}$ with |$\absolutevalue{f^{\prime}(a)-f(a)}\geq\varepsilon$ cannot possibly be minimizing $\mathtt{MSE}_{\tau}(\cdot)$. That is to say, the reward vector $f_{\tau}$ minimizing $\mathtt{MSE}_{\tau}(\cdot)$ must have |$\absolutevalue{f_{\tau}(a)-f^{*}(a)}<\varepsilon$. Then, by the strong notion of $\varepsilon$-self-identifiability, we precisely know that arm $a$ is also a suboptimal arm for the reward vector $f_{\tau}$. Hence, we obtain a contradiction, and arm $a$ cannot possibly be chosen this round.
F.2 Failure: Proof of Theorem 6.2(b)
The proof of Theorem 6.2(b) follows the same structure as Theorem 3.3(b) (see Appendix C.2), with a key modification: although we still aim to show that $\mathtt{Greedy}$ becomes permanently stuck on $a_{\mathtt{dec}}$ with constant probability, the regression oracle may no longer return $f_{\mathtt{dec}}$ exactly—its output may fluctuate around $f_{\mathtt{dec}}$ due to reward noise and the continuity of $\mathcal{F}$. Our key insight is that, under suitable probabilistic events, these fluctuations do not change the greedy decision: the regression output may differ slightly from $f_{\mathtt{dec}}$, but the resulting action remains $a_{\mathtt{dec}}$.
Let us define the following events:
$E_{1}=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(a)-(f_{\mathtt{dec}}(a)-\varepsilon/(2\sqrt{K}))}<\varepsilon/(4\sqrt{K})\quad\text{for each arm $a\neq a_{\mathtt{dec}}$}\,\right\},$
$E_{2}=\left\{\,\forall t>T_{0},\;\absolutevalue{\bar{r}_{t}(a_{\mathtt{dec}})-f^{*}(a_{\mathtt{dec}})}\leq\varepsilon/(4\sqrt{K})\,\right\}.$
These events mirror (3.2) and (3.3), but with two changes: (1) we replace the confidence radius $\Gamma(f_{\mathtt{dec}})/2$ by $\varepsilon/(4\sqrt{K})$, and (2) shift the baseline value of $f_{\mathtt{dec}}(a)$ by $-\varepsilon/(2\sqrt{K})$ in $E_{1}$.
We begin with three probabilistic lemmas.
Lemma F.3.
Event $E_{1}$ happens with probability at least $\left[\frac{\varepsilon}{2\sqrt{2\pi\sigma^{2}K}}\exp(-2/\sigma^{2})\right]^{K-1}$.
Proof F.4.
$\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp(-(x-f^{*}(a))^{2}/(2\sigma^{2})).$
For any $x$ in the interval $[f_{\mathtt{dec}}(a)-3\varepsilon/(4\sqrt{K}),f_{\mathtt{dec}}(a)-\varepsilon/(4\sqrt{K})]$, by boundedness of mean reward, we have
|$\absolutevalue{x-f^{*}(a)}\leq 2.$
Then, the density of $x$ at any point on the interval $[f_{\mathtt{dec}}(a)-3\varepsilon/(4\sqrt{K}),f_{\mathtt{dec}}(a)-\varepsilon/(4\sqrt{K})]$ is at least
$\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp(-2/\sigma^{2}).$
Therefore, for any arm $a$, we have the following.
$\Pr[\absolutevalue{\bar{r}_{{\mathtt{warm}}}(a)-(f_{\mathtt{dec}}(a)-\varepsilon/(2\sqrt{K})}<\varepsilon/(4\sqrt{K})]$ $\geq\frac{\varepsilon}{2\sqrt{2\pi\sigma^{2}K}}\exp(-2/\sigma^{2}).$
$\left[\frac{\varepsilon}{2\sqrt{2\pi\sigma^{2}K}}\exp(-2/\sigma^{2})\right]^{K-1}.$
Lemma F.5.
For some appropriately chosen $\sigma=\Theta(\varepsilon/\sqrt{K})$, we have event $E_{2}$ happens with probability at least
$\Pr{E_{2}}\geq 0.9.$
Proof F.6.
Denote the bad event
$E_{3}=\left\{\exists t>T_{0},\absolutevalue{\bar{r}_{t}-f_{\mathtt{dec}}(a_{\mathtt{dec}})}>\varepsilon/(4\sqrt{K})\right\},$
which is the complement of $E_{2}$. We will obtain an upper bound on $E_{3}$, therefore a lower bound on $E_{2}$. Note that event $E_{2}$ (and $E_{3}$) is only about the decoy arm $a_{\mathtt{dec}}$, and recall that $f^{*}(a_{\mathtt{dec}})=f_{\mathtt{dec}}(a_{\mathtt{dec}})$ by the definition of a decoy.
By union bound,
$\Pr[E_{3}]$ $\leq\sum_{t=1}^{T}\Pr[\absolutevalue{\bar{r}_{t}-f_{\mathtt{dec}}(a_{\mathtt{dec}})}>\varepsilon/(4\sqrt{K})]$
$\leq 2\sum_{t=1}^{T}\exp(-t\varepsilon^{2}/(4\sigma^{2}K))$
$\leq 2\frac{\exp(-\varepsilon^{2}/(4\sigma^{2}K))}{(1-\exp(-\varepsilon^{2}/(4\sigma^{2}K)))}.$
Then, we can choose some suitable $\sigma$ with $\sigma=\Theta(\varepsilon/\sqrt{K})$ ensures $\Pr[E_{3}]<0.1$ and that $\Pr[E_{2}]>0.9$.
Lemma F.7.
For some appropriately chosen $\sigma=\Theta(\varepsilon/\sqrt{K})$, we have the following lower bound:
Proof F.8.
Since event $E_{1}$ concerns all $a\neq a_{\mathtt{dec}}$ and event $E_{2}$ concerns $a_{\mathtt{dec}}$, we know that events $E_{1}$ and $E_{2}$ are independent. As a result,
$\Pr[E_{1}\cap E_{2}]$ $=\Pr[E_{1}]\Pr[E_{2}]$
$\geq 0.9\Pr[E_{1}]$
$=\Omega\left(\left[\frac{\varepsilon}{2\sqrt{2\pi(\varepsilon^{2}/K)K}}\exp(-2/\sigma^{2})\right]^{K-1}\right)$
$=\left[\Omega(\exp(-2/\sigma^{2}))\right]^{K-1},$
where we utilize the previous two lemmas.
Having obtained the previous three probabilistic lemmas, we now prove a crucial lemma which is an extension of Lemma C.7 to the infinite $\mathcal{F}$ setting. At this point, the key insight introduced in Section 6 plays a central role in the proof.
Lemma F.9.
Proof F.10.
The proof is by induction. Assume by round $t$, the algorithm have only choose the decoy arm $a_{\mathtt{dec}}$. Note that assuming event $E_{1}$ and $E_{2}$ holds. Consider the reward function $f_{t}^{\text{emp}}$ given by the empirical means: $f_{t}^{\text{emp}}(a)=\bar{r}_{t}(a)$ for all $a\in\mathcal{A}$. By the induction assumption, $\bar{r}_{t}(a)=\bar{r}_{\text{warm}}(a)$ for each arm $a\neq a_{\mathtt{dec}}$. Hence, by the definition of $E_{1}$ and $E_{2}$, we have
$\|f_{t}^{\text{emp}}-f_{\mathtt{dec}}\|_{2}\leq\sqrt{\sum_{a\in\mathcal{A}}\left(\frac{3\varepsilon}{4\sqrt{K}}\right)^{2}}=\frac{3\varepsilon}{4}.$
Since $f_{\mathtt{dec}}$ is an $\varepsilon$-interior with respect to $\mathcal{F}$, we have $f_{t}^{\text{emp}}\in\mathcal{F}$.
Clearly, $f_{t}^{\text{emp}}\in\mathcal{F}$ is the unique minimizer of $\mathtt{MSE}_{t}(\cdot)$. To see this, for any reward vector $f\neq f_{t}^{\text{emp}}$, we will have
|$\absolutevalue{\bar{r}_{t}(a)-f_{t}^{\text{emp}}}=0\leq\absolutevalue{\bar{r}_{t}(a)-f(a)},$
with at least one inequality strict for one arm. Hence the regression oracle will choose $f_{t}=f_{t}^{\text{emp}}$.
Although $f_{t}^{\text{emp}}$ is not the same as $f_{\mathtt{dec}}$, its optimal action is $a_{\mathtt{dec}}$ when $E_{1}$ and $E_{2}$ happen. This is because
$f_{t}^{\text{emp}}(a_{\mathtt{dec}})$ $\geq f^{*}(a_{\mathtt{dec}})-\frac{\varepsilon}{4\sqrt{K}}$
$=f_{\mathtt{dec}}(a_{\mathtt{dec}})-\frac{\varepsilon}{4\sqrt{K}}$
$=(f_{\mathtt{dec}}(a_{\mathtt{dec}})-\frac{\varepsilon}{2\sqrt{K}})+\frac{\varepsilon}{4\sqrt{K}}$
$>f_{t}^{\text{emp}}(a)$
for all $a\neq a_{\mathtt{dec}}$, where the first inequality follows from the definition of $E_{2}$, the first equality follows from the definition of a decoy, and the last inequality follows from the definitions of $E_{1}$ and $f_{t}^{\text{emp}}$.
Theorem 6.2(b) directly follows from the above lemmas.